OpenStudy (anonymous):

Help me simplify -->13+5(m-4)

OpenStudy (anonymous):

13+5m-20 -13 -13 5m = -33 /5 /5 m = -33/5 -33/5 m = -6.6

OpenStudy (anonymous):

13+5M-20=5M-7